The estimated average concentration of NO2 in air is 0.016 ppm. (a) Calculate the partial pressure of the NO2 in a sample of this air when the atmospheric pressure is 755 torr (99.1 kPa). (b) How many molecules of NO2 are present under these conditions at 20 °C in a room that measures 15 x 14 x 8 ft?

1
Expert's answer
2021-12-24T04:05:09-0500

(a)C=0.016ppm

"P(NO2)=(0.016\/10^6)\u00d7761" =0.00001218

=1.2×10^5torr

b)"PV= nRT"

T= 293.15K

R=8.314

V= 15×14×8 =47.6m³

n=((1.2×10^5)×47.6)/(293.15×8.314)

n= 2.3×10-7moles